A solution contains 0.525 M NaCH_3COO and 0.465 M
Answer #1

For acetic acid, Ka = 1.75 \times 10^{-5}
pKa = -log Ka = - log 1.75 \times 10^{-5} = 4.76
Initially, number of millimoles of sodium acetate = 50.0 \times 0.525 = 26.25
Number of millimoles of acetic acid = 50.0 \times 0.465 = 23.25
15.0 mL of 0.2 M NaOH is added.
Number of millimoles of NaOH = 15.0 \times 0.2 = 3.0
3.0 mmol NaOH will neutralize 3.0 mmol acetic acid to form 3.0 mmol sodium acetate.
Total number of mmol of sodium acetate = 26.25 + 3.0 = 29.25
Total number of mmol of acetic acid = 23.25-3.0 = 20.25

pH = pKa + log \dfrac {[CH_3COONa]}{[CH_3COOH]} \\ pH = 4.76 + log \dfrac {29.25}{20.25} \\ pH = 4.92